# Quantum memory at nonzero temperature in a thermodynamically trivial system

### v1

@@ A.3 Lower bound on the classical mixing time
We also emphasize that Corollary A.13 implies that the thermal Gibbs sampler can be justifiably called a thermal/passive memory. While it is true that a typical system will not be found in a codeword state after finite time due to a nonzero density of small errors, it is the case that one can use standard decoders such as bit flip Sipser and Spielman (1996) or belief propagation [^1] While belief propagation decoding lacks a proof of correctness due to the presence of loops in expander graphs, it has excellent performance in practice Mackay and Neal (1996). Kschischang et al. (2001) to decode the LDPC code at any time before the hitting time of Theorem A.12 and obtain the exact ground state (codeword). From the error correction perspective therefore, these thermal systems serve as excellent memories of the stored information, which can be efficiently decoded.
@@ B.1 CSS codes and the hypergraph product
The easiness of obtaining good classical LDPC codes ironically causes a considerable challenge for their quantum counterparts. Because a random classical LDPC code has a large distance with high probability (recall Appendix A.1), choosing a random LDPC code for $\mathcal{C}_{X}$ will typically result in any orthogonal $\mathcal{C}_{Z}$ to have large weight, and thus be non-LDPC.
The hypergraph product (HGP) was the first quantum LDPC construction to achieve codes with constant rate $K=\Theta(N)$ while maintaining a polynomial scaling of the code distance $D=\Theta(\sqrt{N})$ Tillich and Zemor (2014). The construction takes in two arbitrary classical input codes, described by the 3-term chain (18), and takes their homological tensor product to form a 5-term chain, from which the corresponding CSS code can be identified using (45). The $X$ and $Z$-type parity-check matrices are given by
@@
from which one can straightforwardly verify $H_{X}H^{\mathsf{T}}_{Z}=2H_{1}\otimes H^{\mathsf{T}}_{2}=0$ over $\mathbb{F}_{2}$. Geometrically, the Tanner graph of the HGP code resembles a Euclidean graph product of those of its input classical codes (see Fig. 2). The quantum code parameters are given by
@@
The relative simplicity of the HGP construction leads to the quantum code inheriting many properties of its classical input codes. For ease of notation, we use the same input code (twice) to form the HGP code. If the classical input code is a $(\Delta_{B},\Delta_{C})$ LDPC code, then the associated HGP code is $(\max(\Delta_{B},\Delta_{C}),\Delta_{B}+\Delta_{C})$ LDPC: each physical qubit participates in at most $\max(\Delta_{B},\Delta_{C})$ $X$-checks and likewise $Z$-checks, and each check acts on at most $\Delta_{B}+\Delta_{C}$ physical qubits. If the input classical code is an expander code, then the resulting HGP code is called a quantum expander code. We define the $(\Delta_{B},\Delta_{C})$-HGP ensemble as the collection of CSS codes produced from the hypergraph product of the $(\Delta_{B},\Delta_{C})$-LDPC ensemble.
@@ B.3 Lower bound on the quantum mixing time
We now introduce a definition of non-decodable state, which is a state that could possibly trick a maximum-likelihood decoder, in the worst case. In particular, the following definition will ensure that if we start with a codeword, any logical error necessarily will have passed through a non-decodable state.
@@
We define a non-decodable state to be one with at least one $(s,\ell)$-cluster with size $s\geq D$ and $\ell\geq s/2$ errors.
We are now ready to bound the mixing time of the Gibbs sampler. The idea will be to use the linear confinement property (Corollary B.3) to show that the rate of producing large non-decodable clusters is exponentially suppressed below a critical temperature.
@@
Let’s determine what we should choose for our bottleneck set $E$. From Def. B.7, we know that a non-decodable state must contain at least one $(s\geq D,\ell\geq s/2)$-cluster. We accordingly choose our bottleneck set $E$ be the cut set of states that contain exactly one $(s=\gamma m,\ell\geq fs)$-cluster and no larger, with $f\leq\frac{1}{2}(1-1/s)$. Note that any local trajectory through state space to a non-decodable state must pass through some $\mathbf{x}_{t}\in E$ at some time $t$. The “no larger” condition also automatically implies that our cluster has an error-free perimeter, since a $(s,\ell)$-cluster with an error on its perimeter can always be enlarged to a $(s+1,\ell+1)$-cluster.
@@
$\leq\frac{\pi[E]}{\pi[A]}\leq N\times\exp\left[-\gamma m\left(\frac{2}{3}\beta f-\log(2z)\right)\right]\,.$ (59)
In the second line we used detailed balance (41) and summed over states in $A$; in the third line we bounded $\mathbf{P}_{t}[\mathbf{x}\rightarrow A]\leq 1$ and used (B.3). Inverting the above transition probability gives us a lower bound on the expected hitting time:
@@
Thus, the hitting time will be exponentially long in the polynomial system size $m=\Theta\big{(}\sqrt{N}\big{)}$ as long as
$\beta>\frac{3}{2f}\log(2z)=3\left(1+\frac{1}{\gamma m-1}\right)\log(2z)\,.$ (61)
@@ Appendix C Numerical simulations
In this appendix, we conduct numerical simulations of passive error correction on both the classical and quantum LDPC codes previously discussed. For the classical codes, we randomly sample from the $(4,5)$-LDPC ensemble using the configuration model: we initialize $n$ bits and $\frac{5}{4}n$ checks and pair up their emanating edges according to a random permutation. For the quantum codes, we randomly sample from the $(7,8)$-HGP ensemble by constructing $(7,8)$-LDPC codes using the configuration model and taking the hypergraph product to produce $(8,15)$-LDPC quantum CSS codes. Since the $X$ and $Z$ sectors of the HGP are identical, we focus only on decoding of $X$ errors. All of our sampled codes have no check redundancies. In both classical and quantum simulations, we begin in the $\mathbf{0}$ codeword. We then simulate thermal equilibration at a chosen temperature by performing $\Delta t=100$ Monte Carlo sweeps of the Metropolis algorithm, where each sweep updates all sites exactly once. After the Monte Carlo sweeps, we feed the error syndrome into a belief-propagation decoder Kschischang et al. (2001) and verify if the output codeword is $\mathbf{0}$. For HGP codes, the presence of small loops in their Tanner graphs prevents the convergence of belief propagation (BP), and so we also perform ordered-statistics post-processing Panteleev and Kalachev (2021). We use open source software to implement belief propagation Roffe (2022) and ordered-statistics decoding (BP+OSD) Roffe et al. (2020). In particular, we perform a maximum of 10 iterations of minimum-sum BP with the additional use of OSD-0 for quantum decoding. We sweep from low to high temperatures and terminate the simulation if we fail to decode all Monte Carlo samples at a given temperature. For comparison, we also simulate thermal equilibration for the 2D surface code with minimum-weight perfect-matching (MWPM) decoding, using the open-source software PyMatching Higgott and Gidney (2023).
@@
Appendix D Open quantum system dynamics

### v2

@@ A.3 Lower bound on the classical mixing time
We also emphasize that Corollary A.13 implies that the thermal Gibbs sampler can be justifiably called a thermal/passive memory. While it is true that a typical system will not be found in a codeword state after finite time due to a nonzero density of small errors, it is the case that one can use standard decoders such as bit flip Sipser and Spielman (1996) or message passing Kschischang et al. (2001) to decode the LDPC code at any time before the hitting time of Theorem A.12 and obtain the exact ground state (codeword). From the error correction perspective therefore, these thermal systems serve as excellent memories of the stored information, which can be efficiently decoded.
@@ B.1 CSS codes and the hypergraph product
The easiness of obtaining good classical LDPC codes ironically causes a considerable challenge for their quantum counterparts. Because a random classical LDPC code has a large distance with high probability (recall Section A.1), choosing a random LDPC code for $\mathcal{C}_{X}$ will typically result in any orthogonal $\mathcal{C}_{Z}$ to have large weight, and thus be non-LDPC.
The hypergraph product (HGP) was the first quantum LDPC construction to achieve codes with constant rate $K=\Theta(N)$ while maintaining a polynomial scaling of the code distance $D=\Theta(\sqrt{N})$ Tillich and Zemor (2014). The construction takes in two arbitrary classical input codes, described by the 3-term chain (14), and performs a homological tensor product to form a 5-term chain, from which the corresponding CSS code can be identified using (41). Specifically, the $X$ and $Z$-type parity-check matrices are given by
@@
from which one can straightforwardly verify $H_{X}H^{\mathsf{T}}_{Z}=2H_{1}\otimes H^{\mathsf{T}}_{2}=0$ over $\mathbb{F}_{2}$. Geometrically, the Tanner graph of the HGP code resembles a Euclidean graph product of those of its input classical codes (see Fig. 2 of the main text). The quantum code parameters are given by
@@
The relative simplicity of the HGP construction leads to the quantum code inheriting many properties of its classical input codes. For ease of notation, we use the same input code (twice) to form the HGP code. If the classical input code is a $(\Delta_{B},\Delta_{C})$ LDPC code, then the associated HGP code is $(\max(\Delta_{B},\Delta_{C}),\Delta_{B}+\Delta_{C})$ LDPC: each physical qubit participates in at most $\max(\Delta_{B},\Delta_{C})$ $X$-checks and likewise $Z$-checks, and each check acts on at most $\Delta_{B}+\Delta_{C}$ physical qubits. If the input classical code is an expander code, then the resulting HGP code is called a quantum expander code Leverrier et al. (2015). We define the $(\Delta_{B},\Delta_{C})$-HGP ensemble as the collection of CSS codes produced from the hypergraph product of the $(\Delta_{B},\Delta_{C})$-LDPC ensemble.
@@ B.3 Lower bound on the quantum mixing time
We now introduce a definition of non-decodable state, which is a state that could possibly trick a maximum-likelihood decoder in the worst case. In particular, the following definition will ensure that if we start with a codeword, any logical error necessarily will have passed through a non-decodable state.
@@
A non-decodable state is a state with at least one $(s,\ell)$-cluster with size $s\geq D$ and $\ell\geq s/2$ errors.
We are now ready to bound the mixing time of the Gibbs sampler. The idea will be to use the linear confinement property (Corollary B.3) to show that the rate of producing large non-decodable clusters is exponentially suppressed in the cluster size below a critical temperature.
@@
Let’s determine what we should choose for our bottleneck set $E$. From Def. B.7, we know that a non-decodable state must contain at least one $(s\geq D,\ell\geq s/2)$-cluster. We accordingly choose our bottleneck set $E$ be the cut set of states that contain exactly one $(s=\gamma m,\ell\geq fs)$-cluster and no larger, with $f\leq\frac{1}{2}(1-2/s)$. Note that any local trajectory through state space to a non-decodable state must pass through some $\mathbf{x}_{t}\in E$ at some time $t$. The “no larger” condition also automatically implies that our cluster has an error-free perimeter, since a $(s,\ell)$-cluster with an error on its perimeter can always be enlarged to a $(s+1,\ell+1)$-cluster.
@@
$\leq\frac{\pi[E]}{\pi[A]}\leq N\times\exp\left[-\gamma m\left(\frac{2}{3}\beta f-\log(2z)\right)\right]\,,$ (55)
where in the second line we used detailed balance (37) and summed over states in $A$; in the third line we bounded $\mathbf{P}_{t}[\mathbf{x}\rightarrow A]\leq 1$ and used (B.3). Inverting the above transition probability gives us a lower bound on the expected hitting time:
@@
Thus, the hitting time to a non-decodable state will be exponentially long in the polynomial system size $m=\Theta\big{(}\sqrt{N}\big{)}$ as long as
$\beta>\frac{3}{2f}\log(2z)>3\left(1+\frac{2}{\gamma m-2}\right)\log(2z)\,.$ (57)
@@ Appendix C Numerical simulations
In this section, we conduct numerical simulations of passive error correction on both the classical and quantum LDPC codes previously discussed. For the classical codes, we randomly sample from the $(4,5)$-LDPC ensemble using the configuration model: we initialize $n$ bits and $\frac{5}{4}n$ checks and pair up their emanating edges according to a random permutation. For the quantum codes, we randomly sample from the $(7,8)$-HGP ensemble by constructing $(7,8)$-LDPC codes using the configuration model and taking the hypergraph product to produce $(8,15)$-LDPC quantum CSS codes. Since the $X$ and $Z$ sectors of the HGP are identical, we focus only on decoding of $X$ errors. All of our sampled codes have no check redundancies. In both classical and quantum simulations, we begin in the $\mathbf{0}$ codeword. We then simulate thermal equilibration at a chosen temperature by performing $\Delta t=100$ Monte Carlo sweeps of the Metropolis algorithm, where each sweep updates all sites exactly once. After the Monte Carlo sweeps, we feed the error syndrome into a belief-propagation decoder Kschischang et al. (2001) and verify if the output codeword is $\mathbf{0}$. For HGP codes, the presence of small loops in their Tanner graphs prevents the convergence of belief propagation (BP), and so we also perform ordered-statistics post-processing Panteleev and Kalachev (2021). We use open source software to implement belief propagation Roffe (2022) and ordered-statistics decoding (BP+OSD) Roffe et al. (2020). In particular, we perform a maximum of 10 iterations of minimum-sum BP with the additional use of OSD-0 for quantum decoding. We sweep from low to high temperatures and terminate the simulation if we fail to decode all Monte Carlo samples at a given temperature. For comparison, we also simulate thermal equilibration for the 2D surface code with minimum-weight perfect-matching (MWPM) decoding, using the open-source software PyMatching Higgott and Gidney (2023).
@@
Appendix D Codes built from tree graphs
D.1 Repetition code on a classical tree
In this subsection, we review why the mixing times of classical Ising models on trees are exponentially long in the levels of the trees. On a general graph $G=(V,E)$, the Ising model Hamiltonian is
=$\mathcal{H}=-\sum_{\{i,j\}\in E}Z_{i}Z_{j}.$ (61)
Using (2) in the main text, one can convert this to a parity-check matrix $H$ with row weight $\Delta_{C}=2$. It is straightforward to see that on a connected graph this is simply a repetition code, with all bits in the same state. If the graph has no loops, then there are no redundant parity checks, and Theorem B.4 implies that the system is thermodynamically trivial.
Despite this, we first show that there is nontrivial dynamics, at least on a certain tree:
Theorem D.1.
Consider a $r$-level tree with $n=1+3+\cdots+3^{r}=\frac{1}{2}(3^{r+1}-1)$ vertices, where each interior vertex has exactly 3 children. Then for $\beta$ larger than O(1) constant, the mixing time $t_{\mathrm{mix}}\geq\exp[\mathrm{O}(r)]=n^{\mathrm{\Omega}(1)}$ under Glauber dynamics. Hence this system has thermal self-correction.
This result was proved rigorously in Berger et al. (2005), but we will find it helpful for our later discussion to present a slightly modified proof which follows our proof of Theorem B.8 more closely.
Proof.
Given a state of the spins $\mathbf{z}=\pm 1$ on the tree, define the recursive majority function $g:\mathbb{Z}^{n}_{2}\rightarrow\mathbb{Z}^{n}_{2}$ as follows. For a given state, we label all the boundary vertices by the value of their spins. Next, we label their parent vertices $w$ with whatever state the majority of the children of $w$ are in. We inductively label all the interior vertices from the labels of their children and the value of the recursive majority function $g$ is the label of the root.
Let $S_{+}$ and $S_{-}$ be the set of states with $g=1$ and $-1$. We choose the bottleneck set $\partial S_{-}\subset S_{+}$ as states that are one flip away from $S_{-}$. Since the function $g$ only depends on boundary vertices, for each state in $\partial S_{-}$, we can move the state to $S_{-}$ by flipping one vertex $v$ on the boundary (note that $v$ may not be unique). If we connect any such $v$ and the root with a path, for each interior vertex on the path, their label is +1 and there is one and only one child with label -1. We call such a path the fault path and denote it by $\varphi$.
For a given configuration $\mathbf{z}$, we denote the label of any vertex $v$ by $m_{v}=g_{v}(\mathbf{z})$. Let $F^{-}_{\varphi}$ be the set of states in $\partial S_{-}$ with a specific fault path $\varphi$ and $W^{-}_{\varphi}$ be the set of states in $F^{-}_{\varphi}$ where all the vertices on the fault path have the same label as their spin, +1. We first study only the configurations in $W^{-}_{\varphi}$. Letting $\pi\propto\exp[-\beta\mathcal{H}]$ denote the Gibbs probability distribution, we have:
Figure 5: The two steps of the mapping between $F^{-}_{\varphi}\in S_{-}$ and its complement $\bar{F}^{-}_{\varphi}\in S_{+}$ are depicted. (a) The first step involves flipping the branch of any -1 spin (shaded gray) on the fault path, thereby mapping $F^{-}_{\varphi}\rightarrow W^{-}_{\varphi}$. (b) The second step involves the inductive strategy (green circles) described in Lemma D.2.
Lemma D.2.
If $\beta>\log 2$, the equilibrium probability of states in $W^{-}_{\varphi}$ is exponentially small in $r$:
$\pi[W^{-}_{\varphi}]<\mathrm{e}^{-2(r-1)(\beta-\log 2)}.$ (62)
Proof.
The strategy will follow closely the proof of Theorem B.8: we aim to define a “partner” set $\bar{W}^{-}_{\varphi}\subset S_{+}$, chosen so that we can compare the equilibrium probability of a state $\mathbf{z}\in W^{-}_{\varphi}$ as
$\frac{\pi[\mathbf{z}\in W^{-}_{\varphi}]}{\pi[\bar{\mathbf{z}}\in\bar{W}^{-}_{\varphi}]}=\mathrm{e}^{-\beta[E(\mathbf{z})-E(\bar{\mathbf{z}})]}\,,$ (63)
where $\bar{\mathbf{z}}\in\bar{W}^{-}_{\varphi}$ is the “partner” of $\mathbf{z}$. Intuitively, the partner should have significantly fewer domain walls, so that the ratio in (63) is as small as possible; also, not too many $\mathbf{z}$ should share a partner $\bar{\mathbf{z}}$, so that we can eventually bound $\pi(W^{-}_{\varphi})/\pi(\bar{W}^{-}_{\varphi})$. Define a mapping $W^{-}_{\varphi}\rightarrow\bar{W}^{-}_{\varphi}$ as follows: for each vertex $v$ on the fault path, we inductively apply the following strategy to $v$, its children, grandchildren, etc., until we reach the boundary. For a given vertex, we check whether its children whose labels are $m=-1$ also have spin $Z=-1$. For every such child with both label and spin being -1, we flip the whole branch starting from the child. For every child with label -1 but spin 1, we apply the above strategy to its child: see Fig. 5(b). After the mapping, the energy will always be lower. Note that for states $\mathbf{z}\in W^{-}_{\varphi}$, partnering is not a one-to-one mapping: multiple $\mathbf{z}$ can share the same $\bar{\mathbf{z}}$.
As a simple example, consider states where all the children of vertices on the fault path with label -1 also have spin -1: label this special subset $Y^{-}_{\varphi_{0}}\subset W^{-}_{\varphi}$. To construct the partner subset $\bar{Y}^{-}_{\varphi_{0}}\subset\bar{W}^{-}_{\varphi}$, we will flip all the branches starting from these children, and the energy is lowered by $2(r-1)$. By construction, for any vertex $v$ on the fault path, 2 of the children of $v$ have $m=+1$ and 1 has $m=-1$. Given a known fault path $\varphi_{0}$, there are then at most $2^{r-1}$ possible $\mathbf{z}$ that can map to a given $\bar{\mathbf{z}}$, where $2^{r-1}$ simply counts the possible configurations of the -1 children in the tree: $v$ has 3 children, one of which lies along the known fault path and therefore necessarily has $m=1$). Using (63) we deduce
$\pi[Y^{-}_{\varphi_{0}}]\leq 2^{r-1}\pi[\bar{Y}^{-}_{\varphi_{0}}]\mathrm{e}^{-2(r-1)\beta}\leq\mathrm{e}^{-(2\beta-\log 2)(r-1)},$ (64)
showing that once $\beta>\frac{1}{2}\log 2$, $\pi[Y^{-}_{\varphi_{0}}]$ is very small for thermodynamically large trees.
We now must consider all the possible ways to build $\bar{\mathbf{z}}$ for more general fault paths, weighted by the Boltzmann penalty factor as in (63). For a specific vertex $v$ on the fault path, if its child with label -1 has spin -1, the energy will be lowered by 2 by flipping that child and all of its grandchildren. Following the same steps as above, we deduce a relative Boltzmann weight of
[…]
Next, we consider the relative weight coming from vertices on the fault path where the child with label -1 has spin 1 but all the grandchildren with label -1 have spin -1. Again, there are 2 choices of children that could have label $-1$, but now we know that either 2 or 3 of the relevant grandchildren have spin -1 and label -1. We will flip all of these children. The relative weight coming from these grandchildren flips then
$2c_{2}(\beta)-2c_{1}(\beta):=2\left(3\mathrm{e}^{-4\beta}+\mathrm{e}^{-6\beta}\right),$ (66)
with the first term coming from 2 grandchildren with spin -1, and the second term coming from 3. We’ve defined $c_{2}(\beta)$ such that we can now adjust (65) by replacing $c_{1}$ with $c_{2}$. We can continue to iterate this argument for a given vertex: at the next level, we find
$2c_{n+1}(\beta)-2c_{1}(\beta)=2\left(3c_{n}(\beta)^{2}+c_{n}(\beta)^{3}\right).$ (67)
If $\beta\geq 2\log 2$, it can be checked that $c_{2}(\beta)<2\mathrm{e}^{-2\beta}<1$ and $c_{3}(\beta)<\mathrm{e}^{-2\beta}+4c_{2}(\beta)^{2}\leq 2\mathrm{e}^{-2\beta}$; repeating this argument we deduce that for any $n$,
$c_{n}(\beta)<2\mathrm{e}^{-2\beta}.$ (68)
Now we see that we can just replace $c_{1}$ with $2\mathrm{e}^{-2\beta}$ in (65), which leads to (62), after summing over $\bar{\mathbf{z}}$ and using $\pi[\bar{W}^{-}_{\varphi}]<1$. ∎
Now we return to the more general case:
Lemma D.3.
If $\beta\geq\frac{3}{2}\log 2$, the equilibrium probability of states in $F^{-}_{\varphi}$ is exponentially small in $r$:
$\pi[F^{-}_{\varphi}]<\mathrm{e}^{-(2\beta-3\log 2)(r-1)}.$ (69)
Proof.
For states in $F^{-}_{\varphi}\backslash W^{-}_{\varphi}$, there are some interior vertices $v$ on the fault path with spin -1. For each such $v$, consider the operation $\tilde{g}$ wherein we flip $v$ and all the branches starting from the children of $v$ that are not on the fault path, see Fig. 5(a). In this way, only parity checks that are on the fault path can be changed. After the flipping, there are no violated parity checks on the fault path, while the parity of all the checks off the fault path is unchanged by construction. Hence, $\tilde{g}$ always lowers the energy of the state.
Since each state on the fault path has either $z=1$ or $z=-1$, we conclude that at most $12^{r-1}-$ 1 states in $F^{-}_{\varphi}\backslash W^{-}_{\varphi}$ that can be mapped to the same state in $W^{-}_{\varphi}$ by $\tilde{g}$. Notice that $\tilde{g}$ does not change the label of any vertex on the fault path: because each flip induces by $\tilde{g}$ flips 2 whole branches of children, one of which came with label +1 and the other with label -1, the flip simply exchanges the ±$\pm 1$ label children, while leaving the label of the fault path vertex unchanged. Therefore, $\pi[F^{-}_{\varphi}]<2^{r-1}\pi[W^{-}_{\varphi}]$, and employing (62) we obtain (69). ∎
We now complete the proof of the theorem. For $\partial S_{-}$, there are $3^{r-1}$ different fault paths. Therefore,
$\pi[\partial S_{-}]<3^{r-1}\pi[F^{-}_{\varphi}]<\mathrm{e}^{-(2\beta-\log 24)(r-1)}$ (70)
and the mixing time
$t_{\text{mix}}\geq\frac{\pi[S_{-}]}{4\pi[\partial S_{-}]}>\frac{1}{8}\mathrm{e}^{-(2\beta-\log 24)(r-1)}.$ (71)
We remind the reader that a proof for more general trees is in Berger et al. (2005). ∎
D.2 Hypergraph product codes that are not good quantum memories
For the classical trees, the path $\varphi$ provides a lower bound of the energy barrier that is linear in $r\sim\log n$. We saw that this was, happily, not an impediment to a classical self-correcting memory. However, crucially, the proof of Theorem D.1 was more complicated than Theorem A.12 or Theorem B.8 because we had to be very careful about the possibility of large entropic contributions to the Boltzmann weight of finding clusters of any given size. Put another way, we had to be very explicit to rule out the possibility of an entropy-driven breakdown of thermal memory, even in the presence of the $\mathrm{\Theta}(\log n)$ energy barrier: we did so by showing that a logical error must pass through rare configurations where, at a minimum, there is one flipped parity check at each layer of the tree.
Figure 6: (a) The hypergraph product code of a small degree-3 tree is illustrated. Black dots, blue squares, and red squares represent qubits, $Z$ checks, and $X$ checks respectively. The support of a logical $\bar{X}$ operator is denoted by the qubits inside the shaded red rectangle. (b) A logical $\bar{X}$ operator equivalent to the previous one is drawn.
We now consider the quantum LDPC code obtained from the hypergraph product of the $r$-level classical code with itself. It will still have the same energy barrier, but the probability of the bottleneck set may no longer be assured to be exponentially small in $r$. As shown in Fig. 6, for any given logical operator, by appending appropriate stabilizers that are adjacent to it, we can move the corresponding branch to another row without increasing the energy barrier of the logical operator. In terms of the bottleneck set, we still have $\pi[F^{-}_{\varphi}]\leq\mathrm{e}^{-\mathrm{O}(r)}$, but now the entropic factor is increased by $\lesssim n^{\mathrm{O}(r)}$, where $n$ is the system size of the classical code. Indeed, naively in Fig. 6, we can freely “slide” the support of the logical $\bar{X}$ operator (which runs from left to right) along vertical columns, because the “check-check”-type qubits of the hypergraph product only connect to two $Z$ stabilizers. Suppose we have one of these adversarial error strings that traverse vertically across check-check qubits. If one of its flipped checks is too close to the leaf of the tree, then our large error may be stabilizer-equivalent to smaller errors supported on different rows/trees. However, if the violated parity check is $\ell$ generations removed from the leaves of the tree, it appears to us that there is no such stabilizer reduction to remove a very dangerous entropic factor of at least $3^{\ell}$; if we slide beyond $3^{\ell}$ rows then we again can reduce the weight of the error using stabilizers. We were therefore unable to prove the existence of a bottleneck set whose probability is exponentially small, and of course our above argument suggests in fact that it will not exist – namely, that entropic effects will always dominate. Hence we conjecture that this tree-hypergraph product code is not a passive, self-correcting memory in contrast to the corresponding classical code.
Figure 7: Left: The logical error probability as a function of temperature is plotted for a hypergraph product code consisting of two classical Ising models on 3-regular trees with levels $r_{1}$ (fixed) and $r_{2}$ (variable). Middle: Fixing the logical error probability at 0.25, we interpolate the data on the left plot to estimate the critical inverse temperature as a function of the size of the second classical tree. A linear fit is also overlayed. Right: The logical error probability as a function of temperature is plotted for several codes from the $(4,5)$-HGP ensemble.
We perform numerical simulations of passive error correction for hypergraph product codes of 3-regular trees described above in order to establish confidence for the conjecture; see Fig. 7. To obtain numerical evidence for the entropic breakdown of memory that we argued for above, we fix the size of one tree in the hypergraph product code while varying the number of levels in the other tree, thus making the hypergraph product asymmetric. Explicitly, the $\llbracket N,K,D\rrbracket$ parameters of the tree-hypergraph product code are given by
$N$ $=n_{1}n_{2}+(n_{1}-1)(n_{2}-1)=O\big{(}\mathrm{e}^{r_{1}+r_{2}}\big{)}$ (72a)
$K$ $=1$ (72b)
$D$ $=\min\left(n_{1},n_{2}\right)\,.$ (72c)
The scaling of the critical temperature with the size of the second tree displays a logarithmic dependence which is consistent with free energy arguments: if $c_{1}\sim n\mathrm{e}^{-2\beta}$, for example, we would find that
$\beta_{\mathrm{c}}(n)=\beta_{\mathrm{c},0}+c\log n.$ (73)
To fix the value of $c$, we note that adding long error strings along check-check qubits in the HGP code can be stabilizer-equivalent to a disconnected error cluster, as discussed above. Proposing that we get an extra entropic factor of order $3^{\ell}$ at distance $\ell$ from the leaves, a crude argument suggests that we will gain an added entropic factor of at least $3\times 3^{2}\times 3^{3}\times\cdots\times 3^{r}\sim 3^{r^{2}/2}\sim n^{r/2}$ in (70). This suggests that $c\gtrsim 0.25$, which is actually quite close to what we observe in Fig. 7.
We also numerically simulate several codes from the $(4,5)$-HGP ensemble and do not observe evidence of any decoding threshold (see the rightmost plot in Fig. 7), despite the fact that the classical $(4,5)$-LDPC ensemble is capable of passive error correction (also recall Fig. 4). Our numerics suggest that certain classical passive memories may not have straightforward quantum analogues through the hypergraph product due to the additional degrees of freedom introduced by the graph product. Since it was recently shown that the energy barriers of hypergraph product codes are inherited from their classical input codes Zhao et al. (2024), an interesting direction for future work would be to carefully study the additional entropic effects needed to obtain quantum memory in these codes.
Appendix E Open quantum system dynamics